Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
8x + 11y = -50
-32x – 44y = -200


Answer by mananth(16949)

8x + 11y = -50....................(1)
-32x – 44y = -200.................(2)
when you multiply (1) by 4 we get
-32x+44y=-200
Both the equations are same and they will coincide on the graph
All the points on the line are the solutions to both the equations. So, the system of equations has infinite number of solutions.
So, the system of equations is consistent and the equations are dependent.